OpenStudy (anonymous):

if the parent function is y=1/x describe the changes in the equation y=3/x+1-2.. A) Vertical stretch by a factor of |3|, moves 1 unit to the right, moves down 2 units B)Vertical stretch by a factor of |3|, moves 1 unit to the left, moves down 2 units C) Vertical stretch by a factor of |3|, moves 1 unit to the right, moves up 2 units D) vertical stretch by a factor of |3|, moves 1 unit to the left, moves up 2 units

OpenStudy (jdoe0001):

\(\large { \textit{function transformations} \\ \quad \\ \begin{array}{llll} \begin{array}{llll} shrink\ or\\ expand\\ by\ {\color{purple}{ A}}\end{array} \qquad \begin{array}{llll} vertical\\ shift\\ by \ {\color{green}{ D}} \end{array} \begin{array}{llll}{\color{green}{ D}} > 0& Upwards\\ {\color{green}{ D}} < 0 & Downwards\end{array} \\ % template start \qquad \downarrow\qquad\qquad\quad\ \downarrow\\ y = {\color{purple}{ A}} ( x + {\color{red}{ C}} ) + {\color{green}{ D}}\\ %template end \qquad\qquad \quad \uparrow \\ \qquad\begin{array}{llll} horizontal\\ shift\\ by \ {\color{red}{ C}}\end{array} \begin{array}{llll}{\color{red}{ C}} > 0 & to\ the\ left\\ {\color{red}{ C}} < 0& to\ the\ right\end{array} \end{array} \\ \quad \\ \textit{or if you wish, it's also }\implies y=\cfrac{{\color{purple}{ A}}}{x+{\color{red}{ C}}}+{\color{green}{ D}} }\)

OpenStudy (jdoe0001):

the C component, applies to where the "x" may be encased